# Multi-objective Combinatorial Methodology for Nuclear Reactor Site Assessment: A Case Study for the United States

### v1

@@ 3.1 Non-Dominated, Multi-Objective Combinatorial Search
The objective contribution ($\widehat{OC}$) is row-normalized to not favor any combination length. The indices $i$ and $j$ represent the row and column indices of the matrix, respectively, and are normalized such that the sum of each row equals 1.
$\widehat{NC}_{i,j}=\frac{\widehat{OC}_{i,j}}{\sum_{j}\widehat{OC}_{i,j}}$
@@
The normalized objective contributions ($\widehat{NC}$) are then summed for all combination lengths to get the summed objective contribution ($\widehat{S}$).
$\begin{split}\widehat{S}=\sum_{s=1}^{22}\widehat{NC}\\ \end{split}$
@@
The summed objective contribution ($\widehat{S}$) matrix is row normalized to set the sum of importance to 1 for every site individually in the dataset. The aforementioned operations form the normalized summed objective contribution ($\widehat{SC}$).
$\begin{split}\widehat{SC}_{i,j}=\frac{\widehat{S}_{i,j}}{\sum_{j}\widehat{S}_{i,j}}\end{split}$
@@
An example of achieving the normalized objective contributions is given below. For the selected same objective combination (k=230) with the column numbers 2, 3 and 5, the non-dominated sorting result is given as $\overrightarrow{r}_{230}$. Each row shows a site index which dominates the result of the non-dominated sorting.
$\begin{split}\overrightarrow{r}_{230}&=\begin{bmatrix}0\\ 0\\ 1\\ 1\\ \vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The combination selected for this example includes the columns 2, 3 and 5 of the dataset. The subset including these columns is given as $L_{230}$.
$\begin{split}L_{230}&=\begin{bmatrix}2,3,5\end{bmatrix}\\[10.00002pt] \end{split}$
@@
We use the combination of objectives in this example to create a vector where an entry is 1 if the corresponding objective index is included and 0 otherwise. Note that the length of the vector $\overrightarrow{v_{230}}$ is 22, but only the indices 2, 3, and 5 are non-zero.
$\begin{split}\overrightarrow{v}_{230}&=\begin{bmatrix}0\\ 1\\ 1\\ 0\\ 1\\ \vdots\end{bmatrix}\\[10.00002pt] \end{split}$
The matrix $\overrightarrow{R_{230}}$ is created from the dot product of $\overrightarrow{r}_{230}$ and $\overrightarrow{v}_{230}^{T}$.
$\begin{split}\widehat{R}_{230}&=\begin{bmatrix}0&0&0&0&0&\ldots\\ 0&0&0&0&0&\ldots\\ 0&1&0&1&1&\ldots\\ 0&1&0&1&1&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The contribution matrices of combinations $R_{k}$ are summed for all existing combinations of the given combination length ($s=3$).
$\begin{split}\widehat{OC}&=\begin{bmatrix}0.3&0.5&2&0.7&0.1&\ldots\\ 5&1.1&0.6&0.1&0.3&\ldots\\ 0.2&1.2&0.3&1.7&1&\ldots\\ 0.1&0.2&0.1&1.3&0.5&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The objective contribution matrix of the combination length ($s=3$) is normalized to create the normalized objective contribution matrix $\widehat{NC}$.
$\begin{split}\widehat{NC}&=\begin{bmatrix}0.08&0.14&056&0.19&0.03&\ldots\\ 0.70&0.15&0.08&0.01&0.04&\ldots\\ 0.05&0.27&0.07&0.39&0.23&\ldots\\ 0.05&0.10&0.05&0.55&0.10&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\end{split}$
@@
Figure 2: A multi-objective analysis combinatorial analysis for nuclear reactor sites.
@@
We summarize this method in the flowchart shown in Figure 2. The application of the prescribed method presented several challenges, primarily due to the rapid increase in the number of combinations with varying combination lengths. Prior to the analysis, we executed the siting analysis routine computer core counts and combination lengths. Based on the scaling for core numbers, we anticipated that the complete analysis of all combination lengths would require approximately 8.5 days running on 512 cores. To mitigate the risk of processes being forcibly terminated due to extended runtimes and to address program memory issues, we implemented a checkpoint system between each combination length. We then executed the method in segments, each lasting approximately three days. Each checkpoint records the elapsed computation time for each combination length and the data processed up to that point. We provide a comparison of combination lengths and their corresponding computation times in Table 3.
@@ 3.2 Concatenated Neural Networks (ConcNN)
To evaluate the impact of each parameter on various locations, we propose a Feed-Forward Neural Network (FNN) model. This model aims to predict the siting metric and the importance of site objectives for any location within the contiguous U.S.. Additionally, it seeks to identify which objectives will influence siting decisions, eliminating the need to re-run the high-cost analysis described in the previous section.
@@
The FNN model is composed of two parts. The first part of the model uses the longitude, latitude, county FIPS code (Federal Information Processing Standards), and state FIPS code as its input and the 22 site objectives as the output. The model predicts the siting objectives for a given location by interpolating between the existing locations in the dataset. These predicted objectives are then passed to the second stage of the model. In this stage, the 22 objective predictions, along with the original coordinates, county FIPS, and state FIPS, are input to predict the siting metric (score) and the objective importance values.
@@
To develop the described model, we defined our input data ($X$) with four features, including coordinates, state, and county information. This input passes through a series of hidden layers, producing an intermediate output ($Y_{1}$) that predicts two sets of variables: 14 continuous variables ($Y_{L}$) and 8 binary variables ($Y_{B}$). These variables represent the siting objectives, derived from the STAND tool that are relevant to the chosen location. The intermediate output ($Y_{1}$) is then concatenated with the initial input data ($X$), forming an enhanced feature set. The model then processes this concatenated data through subsequent hidden layers to predict additional siting metrics and objective importance scores ($Y_{2}$). We then train the model to optimize predictions for both $Y_{1}$ and $Y_{2}$ simultaneously. This improved model performance compared to a standard FNN since the initial part of the model is further refined using outputs from the second part.
@@
We conducted hyperparameter tuning using grid search for each of the following parameters: the number of hidden layers of the first part (ranging from 1 to 8), the number of hidden layers of the second part (ranging from 1 to 8), the number of neurons per layer (between 25 and 1,000), and the learning rate (between 1e-5 and 1e-3). The flowchart illustrating the data flow of the model is presented in Figure 3.
To speed up the calculations, we explored an alternative method to replace the first stage of the ConcNN model. This alternative uses a lookup table combined with a linear interpolation scheme to predict NPP siting objectives across locations in the contiguous U.S.. In the original ConcNN model, the first stage processes four input parameters to predict 22 siting objectives. For predicting the characteristics of the locations, the lookup table and interpolation method are expected to perform effectively. This is because most siting objectives are primarily influenced by proximity-based metrics. LUT-NN refers to this revised model, which uses a lookup table for predicting the site objectives. ConcNN will be used to refer to the original, fully neural network model that predicts all outputs.. Later in the results section, we benchmark both LUT-NN and ConcNN in predicting site objectives, the siting metric, and objective importance values. The data flow for both LUT-NN and ConcNN models is given in Figure 3 and 4.
@@
Figure 4: Visualization of the data flow of the LUT-NN model (lookup table plus neural network).
@@ 3.3 Computing Resources
This study used the Idaho National Laboratory High-Performance Computing (INL-HPC) system to conduct the one-time combinatorial search described in Section 3.1. The analysis encompassed all brownfield sites, CPP sites, and the combined dataset of brownfield and CPP sites. The computations were performed in parallel on the Bitterroot cluster, utilizing 512 cores. Each CPU in the cluster is an Intel® Xeon® Platinum 8480+ with 56 cores operating at a frequency of 3.8 GHz, with 256 GB of RAM per node. The search was successfully completed over a span of nine days.
For the machine learning training, we employed TensorFlow 2.18.0 and Keras 3.6.0 to develop and train the ConcNN model, as detailed in Section 3.2. We conducted the training and hyperparameter tuning on an internal GPU server at the University of Michigan. This server is equipped with two AMD EPYC 9654 processors, each providing 96 cores operating at 2.4–3.7 GHz, resulting in a total of 192 cores and 384 threads. Additionally, the server features four NVIDIA RTX 6000 Ada Generation GPUs and 1536 GB of DDR5 RAM. The training and tuning process completed in approximately 18 hours.
@@ 4.1 Reactor Siting Analysis in the United States
We analyzed the brownfield location dataset using the resources explained in Section 3.3 until site observation ratios for all 22 combination lengths were computed based on the method presented in Section 3.1. After this process, we used the resulting data to calculate the siting metric. We used this siting metric to identify optimal locations and evaluate the influence of each objective on the site score. The workflow employed to derive these findings is summarized in Figure 5.
Figure 5: A flowchart summarizing the results acquired from brownfield sites dataset.
@@
To illustrate how the processed data varies with changing combination lengths, we plot the site observation ratios for different combination lengths prior to summing them. The top-performing locations and their observation ratios across various combination lengths are presented in Figure 6.
@@
Figure 6: The normalized observation ratio as a function of combination length from the non-dominated multi-objective combinatorial search. The plot shows the change of $\overrightarrow{NR}$ with combination length ($s$) for the top performing 20 brownfield sites.
@@
Considering the required computational power, we present the non-dominated, multi-objective combinatorial search operation as a brute-force method for this dataset. The number of locations is significantly larger in the brownfield siting dataset (34,212) than the coal site dataset (265). For this reason, non-dominated, multi-objective combinatorial search provides more reliable outcomes on the brownfield NPP siting dataset. We give the product of these operations as a summed and scaled observation ratio (siting metric) for each location. The site scores of the best brownfield sites are given in Table 4.
@@
Table 4: Site scores (summed and scaled observation ratios $\overrightarrow{\mathbb{M}}$) by site ID, state and county for the top five performing brownfield locations in the contiguous U.S..
@@
Brownfield ID State County Longitude Latitude Siting Metric ($\overrightarrow{\mathbb{M}}$)
110038759572 FL St. Lucie -80.320 27.418 1.000
110015334440 CA Almeda -122.103 37.668 0.989
110038756833 MT Garfield -106.907 47.318 0.988
110039543330 CT Hartford -72.674 41.763 0.973
110038712453 NC Wake -78.631 35.763 0.945
@@
For the best-performing site ( ACRES location ID 110038759572, which is located in St. Lucie County, Florida), the change in the observation ratio is illustrated in Figure 7. As shown, the observation ratio tends to decrease with increasing combination length. This trend is consistent across all other sites examined in the procedure outlined in Figure 2. We attribute the observed decrease to the total number of combinations associated with different combination lengths; specifically, fewer combinations at both high and low combination lengths (see Table 2) result in lower observation rates. This becomes more obvious in Figure 8, which shows the change in the normalized observation ratio variance as a function of combination length. Nevertheless, high-order combination lengths (e.g. 18-20) cause larger variance in the observation ratio compared to the low-order ones (e.g., 3-5)
We show one of the most interesting findings in this study in Figures 6-8, which reveal that the most promising sites tend to exhibit a prominent observation rate on the Pareto front at higher combination lengths, peaking between 10-12 combinations. This suggests that the interaction between objectives plays a critical role in site ranking. Simply assigning weights to individual objectives fails to capture these high-dimensional interactions. These higher-order interactions significantly influence site ranking, while lower-order combination lengths have a relatively smaller impact on the variance of observation rate for different sites.
@@
Figure 7: Normalized observation ratio ($\overrightarrow{NR}$) of the best performing brownfield site (ACRES Registry ID 110038759572 in St. Lucie County, Florida) for different combination lengths ($s$).
@@
Figure 8: The variance in the normalized observation ratio as a function of the combination length ($s$). This figure shows how the $\overrightarrow{NR}$ results differ for increasing combination lengths. For $s=1$, the rows of $\overrightarrow{NR}$ are mostly the same (for each site) with approximately 0 variance in the vector.
@@
The top-performing brownfield locations identified in this analysis are situated in St. Lucie County, Florida (FL); Alameda County, California (CA); Garfield County, Montana (MT); Hartford County, Connecticut (CT); and Wake County, North Carolina (NC). Key distinguishing objectives for these sites include distance to retiring facilities, public sentiment toward nuclear energy, proximity to nuclear R&D centers, terrain slopes, and fault line proximity. Of these five sites, only the location in California (2) has a nuclear restriction, which is the need for high-level waste disposal technology or reprocessing capacity. In terms of energy policy and incentives, the top two sites in Florida (1) and California (2) have nuclear-inclusive policies, while the third site in Montana (3) has higher electricity prices. Additionally, the California site (2) operates in a traditionally regulated energy market; all other sites have deregulated markets. California, however, it is the only location adjacent to protected lands. From a safety perspective, only the Montana site (3) is within 200 miles of a fault line, but it ranks positively for all other safety objectives because it is the safest site in the top five results. The California site (2) is the only site with multiple nearby nuclear R&D centers. he Florida site (1) ranks highest for population sentiment favoring nuclear energy.
@@
This analysis shows optimal site selection without weighting the importance of specific objectives,thereby yielding an objective conclusion. This is a stark contrast to the subjectivity of status quo methods. The attributes of the top brownfield sites highlight the balanced nature of competing objectives in the selection process.
To create a joint dataset between brownfield and CPP sites, the top 250 sites from the brownfield dataset were combined with all 265 sites from the CPP dataset. A consistent combinatorial methodology was applied to assess and compare optimal brownfield and coal power plant (CPP) sites, facilitating a comprehensive analysis of the top-performing locations across both types of sites. The best 20 sites from this comparison, along with their coordinates and site metrics, are illustrated in Figure 9.

### v2

@@ 3.1 Non-Dominated, Multi-Objective Combinatorial Search
The objective contribution ($\widehat{OC}$) is row-normalized to not favor any combination length. The indices $i$ and $j$ represent the row and column indices of the matrix, respectively, and column indices ($j$) are normalized such that the sum of each row equals 1.
$\widehat{NC}_{s,i,j}=\frac{\widehat{OC}_{s,i,j}}{\sum_{j}\widehat{OC}_{s,i,j}}\quad,\quad\widehat{NC}\in\mathbb{R}^{22\times 16057\times 22}.$ (7)
@@
The normalized objective contributions ($\widehat{NC}$) are then summed for all combination lengths ($s$) to get the summed objective contribution ($\widehat{S}$).
$\begin{split}\widehat{S}_{i,j}=\sum_{s=1}^{22}\widehat{NC}_{s,i,j}\quad,\quad\widehat{S}\in\mathbb{R}^{16057\times 22}.\\ \end{split}$ (8)
@@
The summed objective contribution ($\widehat{S}$) matrix is row-normalized to set the sum of importances to 1 for every site individually in the dataset. The aforementioned operations form the normalized summed objective contribution ($\widehat{SC}$).
$\begin{split}\widehat{SC}_{i,j}=\frac{\widehat{S}_{i,j}}{\sum_{j}\widehat{S}_{i,j}}\quad,\quad\widehat{SC}\in\mathbb{R}^{16057\times 22}.\end{split}$ (9)
@@
An example of achieving the normalized objective contributions is given below. For the selected same objective combination ($k=230$) with the column numbers 2, 3 and 5, the non-dominated sorting result is given as $\widehat{r}[s=3,k=230]$. Each row shows a site index which dominates the result of the non-dominated sorting.
$\begin{split}\widehat{r}[s=3,k=230]&=\begin{bmatrix}0\\ 0\\ 1\\ 1\\ \vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The combination selected for this example includes the columns 2, 3 and 5 of the dataset. The index set inside $L$ including these columns is given as $L[s=3,k=230]$.
$\begin{split}L_{3,230}&=\begin{bmatrix}2,3,5\end{bmatrix}\\[10.00002pt] \end{split}$
@@
We use the combination of objectives in this example to create a vector where an entry is 1 if the corresponding objective index is included and 0 otherwise. Note that the length of the vector $\overrightarrow{v}[s=3,k=230]$ is 22, but only the indices 2, 3, and 5 are non-zero.
$\begin{split}\widehat{v}[s=3,k=230]&=\begin{bmatrix}0\\ 1\\ 1\\ 0\\ 1\\ \vdots\end{bmatrix}\\[10.00002pt] \end{split}$
The matrix $\widehat{R}_{s,k}$ is created from the dot product of $\widehat{r}[s=3,k=230]$ and $\widehat{v}[s=3,k=230]^{T}$.
$\begin{split}\widehat{R}[s=3,k=230]&=\begin{bmatrix}0&0&0&0&0&\ldots\\ 0&0&0&0&0&\ldots\\ 0&1&0&1&1&\ldots\\ 0&1&0&1&1&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The contribution matrices of combinations $\widehat{R}$ are summed for all existing combinations of the given combination length (e.g., $s=3$).
$\begin{split}\widehat{OC}[s=3]&=\begin{bmatrix}0.3&0.5&2&0.7&0.1&\ldots\\ 5&1.1&0.6&0.1&0.3&\ldots\\ 0.2&1.2&0.3&1.7&1&\ldots\\ 0.1&0.2&0.1&1.3&0.5&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\\[10.00002pt] \end{split}$
@@
The objective contribution matrix of the combination length ($s=3$) is normalized to create the normalized objective contribution matrix $\widehat{NC}$ for $s=3$. Summing all $\widehat{NC}$ results for $s\in[1,22]$ values and row-normalizing this $\widehat{S}$ matrix creates the objective contribution result ($\widehat{SC}$) for each site ($i$) and for each objective ($j$).
$\begin{split}\widehat{NC}[s=3]&=\begin{bmatrix}0.08&0.14&056&0.19&0.03&\ldots\\ 0.70&0.15&0.08&0.01&0.04&\ldots\\ 0.05&0.27&0.07&0.39&0.23&\ldots\\ 0.05&0.10&0.05&0.55&0.10&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{bmatrix}\end{split}$
@@
Figure 2: A flowchart of the multi-objective combinatorial analysis for nuclear reactor sites proposed in this study.
@@
We summarize this method in the flowchart shown in Figure 2. The application of the prescribed method presented several challenges, primarily due to the rapid increase in the number of combinations with varying combination lengths. Prior to the analysis, we executed the siting analysis routine with varying computer core counts and combination lengths. Based on the scaling for core numbers, we anticipated that the complete analysis of all combination lengths would require approximately 5 days running on 360 processors. To mitigate the risk of processes being forcibly terminated due to extended runtimes and to address program memory issues, we implemented a checkpoint system between each combination length. We then executed the method in segments. Each checkpoint records the elapsed computation time for each combination length and the data processed up to that point. We provide a comparison of combination lengths and their corresponding computation times in Table 3.
@@ 3.2 Concatenated Neural Networks (ConcNN)
To evaluate the impact of each parameter on various locations, we propose a Feedforward Neural Network (FNN) model. This model aims to predict the siting metric and the importance of site objectives for any location within the contiguous United States. Additionally, it seeks to identify which objectives will influence siting decisions, eliminating the need to re-run the high-cost analysis described in the previous section. Machine learning applications in nuclear through FNN and other forms of neural networks were broad and include surrogate modeling for expensive reactor simulations [39], digital twins [40], multi-objective optimization for nuclear reactor control [41], fault detection in nuclear power plants [42], time series forecasting of nuclear accidents [43], among others.
@@
The ConcNN model is designed to predict 22 siting objectives, one siting metric, and 22 siting importance values for each site corresponding to each objective. The purpose of creating the neural network model is to decrease the computational time needed for estimating these quantities. This FNN model is composed of two parts. The first part of the model uses the longitude, latitude, county FIPS code (Federal Information Processing Standards [^1] https://transition.fcc.gov/oet/info/maps/census/fips/fips.txt), and state FIPS code as its input and the 22 site objectives as the output. These predicted objectives are then passed to the second stage of the model. In this stage, the 22 objective predictions, along with the original coordinates, county FIPS, and state FIPS, are input to predict the siting metric (score) and the objective importance values. This training approach reflects the hierarchical relationship where the location influences siting objectives, and siting objectives impact the siting metric and the objective importance values. To replicate this dependency within a unified framework, a single model structure, as depicted in Figure 11, is implemented for ConcNN, ensuring end-to-end training and prediction within the same model.
@@
To develop the described model, we defined our input data ($X$) with four features, including coordinates, state, and county information. This input passes through a series of hidden layers, producing an intermediate output ($Y_{1}$) that predicts two sets of variables: 14 continuous variables ($Y_{L}$) and 8 binary variables ($Y_{B}$). These variables represent the siting objectives, derived from the STAND tool that are relevant to the chosen location. The intermediate output ($Y_{1}$) is then concatenated with the initial input data ($X$), forming an enhanced feature set. The model then processes this concatenated data through subsequent hidden layers to predict additional siting metrics and objective importance scores ($Y_{2}$). We then train the model to optimize predictions for both $Y_{1}$ and $Y_{2}$ simultaneously. This could improve model performance compared to a standard FNN since the initial part of the ConcNN model is further refined using outputs from the second part.
@@
We conducted hyperparameter tuning using grid search for each of the following parameters: the number of hidden layers of the first part (ranging from 1 to 8), the number of hidden layers of the second part (ranging from 1 to 8), the number of neurons per layer (between 25 and 1,000), and the learning rate (between 1e-5 and 1e-3). The flowchart illustrating the data flow of the ConcNN model is presented in Figure 3.
@@
3.3 Lookup Table Plus Neural Networks (LUT-NN)
To speed up the calculations and possibly enhance the performance even further, we explored an alternative method to replace the first stage of the ConcNN model. This alternative uses a lookup table combined with a linear interpolation scheme to predict NPP siting objectives across locations in the contiguous United States. In the original ConcNN model, the first stage processes four input parameters to predict 22 siting objectives. For predicting the characteristics of the locations, the lookup table and interpolation method are expected to perform effectively. This is because most siting objectives are primarily influenced by proximity-based metrics, with some depending solely on the state and county in which the site is located. Accordingly, LUT-NN refers to this revised model, which uses a lookup table for predicting the site objectives. ConcNN will be used to refer to the original, fully neural network model that predicts all outputs. Later in the results section, we benchmark both LUT-NN and ConcNN in predicting site objectives, the siting metric, and objective importance values. The data flow schemes for both LUT-NN and ConcNN models are given in Figure 3 and 4, respectively.
A connection can be established between the proposed neural network models and Explainable Artificial Intelligence (XAI) methods. Various approaches to XAI have been explored, while no single method is universally applicable to all machine learning models. However, the performances of XAI techniques have been well-documented for specific methodologies [44, 45]. The research methodology applied in this paper implements a non-dominated multi-objective combinatorial search, designed to aggregate results from each Pareto front where the number of objectives is too large to provide reliable outcomes through traditional non-dominated sorting. This combinatorial search tracks the contribution of each objective to the respective Pareto fronts. By accessing every step of the computation, the methodology ensures transparency, enabling the reconstruction of reasoning for each evaluated combination and demonstrating the influence of individual objectives on the siting metric/score.
In our computation, we sum contributions from different objectives to a single siting metric, mirroring, for example, the popular SHAP (SHapley Additive exPlanations) approach, where the contribution of each feature in the machine learning model is calculated and aggregated to explain the final prediction [46]. Additionally, the step where we normalize the objective contributions aligns with SHAP’s methodology for distributing credit (or contribution) fairly among objectives [47]. SHAP ensures that contributions from all features sum to the predicted outcome (local additivity), similar to our approach. In the context of explainability, our proposed method represents a simplified additive attribution method inspired by SHAP principles.
Figure 4: Visualization of the data flow of the LUT-NN model (Lookup Table Plus Neural Networks).
@@ 3.4 Computing Resources
This study used an internal server at the University of Michigan to evaluate the final results of the joint dataset with Brownfield sites and CPP sites as described in Section 3.1. The combinatorial search was performed using a multithreading calculation of 360 threads (180 cores), and successfully completed over a span of five days. For the machine learning training, we employed TensorFlow 2.18.0 and Keras 3.6.0 to develop and train the ConcNN and LUT-NN models. We conducted the training and hyperparameter tuning on the same internal server, which is equipped with two AMD EPYC 9654 processors, each providing 96 cores operating at 2.4–3.7 GHz, resulting in a total of 192 cores and 384 threads. Additionally, the server features four NVIDIA RTX 6000 Ada Generation GPUs and 1536 GB of DDR5 RAM. The training and tuning processes were completed in approximately 18 hours.
Additionally, the Idaho National Laboratory High-Performance Computing (INL-HPC) system has been used to produce preliminary results during the early stages of this research before moving to the internal server. This transition occurred due to the restrictions of wall time limits that hindered completing our combinatorial search in an effective manner when a large number of sites were used. The computations were performed in parallel on the Bitterroot cluster, utilizing 256 cores. Each CPU in the cluster is an Intel® Xeon® Platinum 8480+ with 56 cores operating at a frequency of 3.8 GHz, with 256 GB of RAM per node.
@@ 4.1 Reactor Siting Analysis in the United States
We analyzed the joint dataset using the resources explained in Section 3.4 until site observation ratios for all 22 combination lengths were computed based on the method presented in Section 3.1. After this process, we used the resulting data to calculate the siting metric/score. We used this siting score to identify optimal locations and evaluate the influence of each objective on the site score. The workflow employed to derive these findings is summarized in Figure 5.
Figure 5: A flowchart summarizing how the site data are combined and processed to feed into the machine learning models
@@
To illustrate how the processed data changes with varying combination lengths, we plot the site observation ratios for different combination lengths prior to summing them. The top-performing locations and their observation ratios across various combination lengths are presented in Figure 6. These results indicate that among the top 10 sites, 6 are coal sites and 4 are Brownfield sites. However, in the top 20 ranking, the numbers shift to 14 coal sites and 6 Brownfield sites.
@@
Figure 6: The normalized observation ratio as a function of combination length from the non-dominated multi-objective combinatorial search. The plot shows the change of $\widehat{NR}$ with combination length ($s$) for the top 20 performing sites.
@@
Considering the required computational power, we present the non-dominated, multi-objective combinatorial search operation as a brute-force method for this dataset. We give the product of these operations as a summed and scaled observation ratio (siting score) for each location. The site scores of the best six sites are given in Table 4, which include four coal sites and two Brownfield sites.
@@
Table 4: Site scores (summed and scaled observation ratios $\overrightarrow{\mathbb{M}}$) by site ID, state and county for the top six locations in the contiguous United States
@@
Site ID State County Longitude Latitude Site Score ($\overrightarrow{\mathbb{M}}$)
C2914 Ohio Tuscarawas -81.47 40.52 1.0000
C2712 North Carolina Person -79.07 36.48 0.7991
C2367 New Hampshire Rockingham -70.78 43.41 0.7025
C8042 North Carolina Stokes -80.06 36.28 0.6650
110038759572 Florida St. Lucie -80.32 27.42 0.6169
110015334440 California Alameda -122.10 37.67 0.6018
@@
For the best-performing site (CPP ID C2914, which is located in Tuscarawas County, Ohio), the change in the observation ratio is illustrated in Figure 7. As shown, the observation ratio tends to decrease with increasing combination length. This trend is consistent across all other sites examined in the procedure outlined in Figure 2. We attribute the observed decrease to the total number of combinations associated with different combination lengths; specifically, fewer combinations at both high and low combination lengths (see Table 2) result in lower observation rates. This becomes more obvious in Figure 8, which shows the change in the normalized observation ratio variance as a function of combination length. Nevertheless, high-order combination lengths (e.g. 18-20) cause larger variance in the observation ratio compared to the low-order ones (e.g., 1-3)
@@
Figure 7: Normalized observation ratio ($\widehat{NR}$) of the best performing CPP site (CPP ID C2914 in Tuscarawas County, Ohio) for different combination lengths ($s$).
We show one of the most interesting findings in this study in Figures 6-8, which reveal that the most promising sites tend to exhibit a prominent observation rate on the Pareto front at middle combination lengths, peaking between 6-9 combinations. This suggests that the interaction between objectives plays a critical role in site ranking. Simply assigning weights to individual objectives fails to capture these high-dimensional interactions. These higher-order interactions significantly influence site ranking, while lower-order combination lengths have a relatively smaller impact on the variance of observation rate for different sites.
As shown in Figure 8, observation ratio variance increases with combination length. For a length of 1, all sites perform nearly identically, but as the length grows, site rankings become more distinct. The highest variance occurs at 7-way combinations, indicating its effectiveness in differentiating sites. For the combination length of 7, the performance disparity is highest between the best and worst sites in the results. For instance, while lower-order combinations (e.g., 1-3) provide limited separation, higher-order ones (e.g., 8-12) introduce greater variability but require significantly more computation time. While assessing combination lengths with the highest variance would significantly reduce computation time, the most impactful length cannot be determined in advance. As a result, this evaluation cannot be limited to the most effective combination lengths, making it necessary to assess all of them for a thorough site ranking.
@@
Figure 8: The variance in the normalized observation ratio as a function of the combination length ($s$). This figure shows how the $\widehat{NR}$ results differ for increasing combination lengths. For $s=1$, the rows of $\widehat{NR}$ are mostly the same (for each site) with approximately 0 variance in the first column of the $\widehat{NR}$ matrix.
@@
The top-performing site locations identified in this analysis based on Table 4 are situated in Tuscarawas County, Ohio (OH); Person County, North Carolina (NC); Rockingham County, New Hampshire (NH); Stokes County, North Carolina (NC), St. Lucie County, Florida (FL); and Alameda County, California (CA). Key distinguishing objectives for these sites include state nuclear inclusive policy, public sentiment toward nuclear energy, number of intersecting protected lands, population center distance, and retiring facility distance. None of these sites has an intersection with a protected land. Of these six sites, only the location in California (6) has a nuclear restriction, which is the need for high-level waste disposal technology or reprocessing capacity. In terms of energy policy and incentives, the sites in Ohio (1) and North Carolina (2, 4) and California (6) have nuclear-inclusive policies, while the site in Florida (5) has highest positive population sentiment towards nuclear energy. The site in Ohio (1) has the lowest substation distance, and has high state electricity imports. The sites in North Carolina (2, 4) has the lowest state construction costs, has all of their safety objectives as positive except the landslide risk, and has no hazardous facilities within 5-miles unlike other sites. Additionally, the top 4 sites have a distance higher than 20 miles to the closest population center. From a safety perspective, the first five sites have most of their safety objectives positive, and are among the safest locations in the whole dataset. The California site (6) is the only site with multiple nearby nuclear R&D centers. However, it is also very close to hazardous facilities and population centers and does not have a streamflow.
@@
This analysis shows optimal site selection without weighting the importance of specific objectives, thereby yielding an objective conclusion. This is a stark contrast to the subjectivity of status quo methods. The attributes of the top Brownfield and CPP sites highlight the balanced nature of competing objectives in the selection process. The best 20 sites from this comparison, along with their coordinates and site metrics, are illustrated in the map in Figure 9.